OpenStudy (anonymous):

1/150 = 6.667x10 ^-3

OpenStudy (anonymous):

If a wave has amplitude of 0.2 meters, a wavelength of 0.5 meters, and a frequency of 32 Hz, at what speed is it moving?

OpenStudy (anonymous):

λ = wavelength = 0.5 m ν = frequency = 32 hz Speed = λ * ν Speed = 0.5 * 32 = 16m/s